mikewwe13:

Select "Growth" or "Decay" to classify each function. Function Growth Decay y=200(0.5)^2t y=1/2(2.5)^t/6 y=(0.65)^t/4 Which equation can be used to find the solution of (1/3)^d−5 = 81 ? A. −d−5=4 B. d + 5 = 4 C. d−5=4 D. −d+5=4

Ultrilliam:

@Vocaloid

Vocaloid:

@mikewwe13 so you determined earlier that (1/3)^(d-5) = (1/3)^-4 so d - 5 = -4 you can multiply both sides of the eqn by -1 to get -d + 5 = 4 (answer D)